# RacerF: Lightweight Static Data Race Detection for C Code

### v1

@@ Abstract
We present a novel static analysis for thread-modular data race detection. Our approach exploits static analysis of sequential program behaviour whose results are generalised for multi-threaded programs using a combination of lightweight under- and over-approximating methods. We have implemented this approach in a new tool called RacerF as a plugin of the Frama-C platform. RacerF can leverage several analysis backends, most notably the Frama-C’s abstract interpreter EVA. Although our methods are mostly heuristic without providing formal guarantees, our experimental evaluation shows that even for intricate programs, RacerF can provide very precise results competitive with more heavy-weight approaches while being faster than them.
@@ 1 Introduction
A lot of work has also been invested into designing various static as well as dynamic methods of finding data races and/or proving their absence. However, the complex behaviour of concurrent programs reflects in that such methods do often have problems with scalability or precision. In this paper, we contribute to the ongoing search for a suitable balance between the scalability and precision of data race detection by proposing a novel static analysis for thread-modular data race detection.
@@
Our approach relies on leveraging results of static analysis of the sequential behaviour of program threads for finding data races in multi-threaded programs. We use two strategies of combining information resulting from analysis of the sequential behaviour of program threads: one under-approximates and one over-approximates the concurrent behaviour of the threads running in parallel. We combine these strategies to obtain a more precise analysis such that the under-approximation is used to find data races while the over-approximation is used to show their absence.
@@
We have implemented our approach in a new data race detector for C programs with POSIX threads (pthreads). The tool is called RacerF and it is implemented as a plugin of the Frama-C platform [9]. RacerF can use several analysis backends implementing the underlying analysis of the sequential behaviour of program threads and differing in their precision and scalability. One of them is built on top of the Frama-C’s value analysis plugin EVA based on abstract interpretation [5], one is purely syntactic, and the last one is based on the Frama-C’s plugin for may-alias and points-to analysis [4].
@@
We have evaluated our implementation on a number of benchmarks coming both from the SV-COMP competition in software verification, which includes both synthetic benchmarks as well as benchmarks derived from real-world code, as well as on a number student projects in an advanced course of operating systems. Although our approach is mostly heuristic without providing formal guarantees, our experimental evaluation shows that even for intricate programs, RacerF can provide very precise results competitive with more heavy-weight approaches while being faster than them.
@@ 2 Related Work
RacerX [13] is a static analyser for C based primarily on a lockset analysis. The tool implements many heuristics to deal with analysis mistakes and has been successfully used to find bugs in the kernels of several operating systems. Compared to our approach, RacerX uses only a very simple pointer analysis for the representation of locks. RacerD [3] is a compositional analyser implemented in the Infer framework. It provides scalability by analysing each function only once. Unlike us, RacerD targets Java programs with structural locking, assuming that all locking operations are balanced. O2 [21] is another static checker building on the notion of origins unifying threads and events. The approach was implemented for C, C++, and Java/Android; and used to find many previously unknown bugs in real-world code. However, as we show in our experiments, it can still miss many data races in smaller but real-world programs.
@@
On the other side, there are tools focused on soundness, rather than race detection. Goblint [26] and Astrée [19, 22] rely on a thread-modular abstract interpretation to soundly over-approximate thread behaviour. The work [26] further proposes a way to handle common locking schemes in Linux driver devices. In our approach, we are not interested in the analysis of multithreaded programs in general and we can consequently afford more drastic approximations that seem to be sufficient for data race detection. We also focus on under-approximation to detect bugs.
@@
As an alternative to static data race detection, various approaches based on testing and dynamic analysis can be used. A big challenge for these approaches is the fact that errors like data races can manifest in very rare behaviours only, which are difficult to catch, even through many repeated test runs. Coverage of such rare behaviours can be improved using approaches such as systematic testing [27] or noise-based testing [11, 15]. Another way is to use extrapolating dynamic checkers, such as [24, 12, 17], which can report warnings about possible errors even if those are not seen in the testing runs, based on spotting their symptoms. However, even though such checkers have proven quite useful in practice, they can, of course, still miss rarely occuring errors. Moreover, monitoring a run of large software through such checkers may be quite expensive too.
@@
Approaches based on model checking, i.e., exhaustive state-space exploration, can guarantee the discovery of all potentially present errors – either in general or at least up to some bound, often given in the number of context switches. However, so far, the scalability of these techniques is not sufficient to handle large industrial code, even when combined with methods such as sequentialisation [20, 23] and others used in the newest tools such as Deagle [14]. Indeed, this shows up in our experiments too.
@@ 3 Preliminaries
We now introduce several basic notions used in the further exposition.
@@
The $\mathsf{CValue}$ domain and memory addresses.
Throughout the paper, we frequently refer to the $\mathsf{CValue}$ abstract domain [5, 9] implemented in Frama-C and used by the abstract interpretation of EVA. We are primarily interested in the representation of memory addresses. A memory address is represented as a pair $(b,o)$ where $b$ is a base address (in the following referred to simply as a base) and $o$ is an offset. Each variable (global, local, and also formal) has its associated base address. Dynamic allocations are modelled by creating so-called dynamic bases which come in two flavours – a strong dynamic base corresponds to the result of a precisely one call to an allocation function; on the other hand, weak memory bases may represent multiple allocated addresses (typically, they are the consequence of allocations in loops). Offsets are represented by a dedicated integer domain, the details of which are not important for this paper. We use $\mathsf{Cvalue.value}$ for the domain representing possible values of a single variable and $\mathsf{Cvalue.state}$ for the domain representing program state, i.e., possible values for multiple variables. We use $\mathsf{Addresses}$ to denote the aforementioned set of memory addresses.
@@
We call an analysis context a triple $(thread,calls,stmt)$ consisting of a thread, a sequence of the latest $n$ calls (each of them formed by a call site and the function called), and a program statement. Our data-race analysis relies on various underlying analyses presented below to compute an abstract representation of thread states reachable by performing the given statement for each analysis context. Notice that a context holds no information about how the thread was created nor how its execution was interleaved with other threads. The empty call sequence is denoted by $\varepsilon$. The number $n$ is a parameter of the analysis set to 1 by default to handle simple lock/unlock wrapper functions. We use $\mathsf{Context}$ to denote the set of all possible contexts.
$\mathsf{analyse\_thr}$ $:\mathsf{Function}\times\mathsf{CValue.state}\xrightarrow{}\mathsf{Result}$ (sequentual analysis of a thread)
@@
$\mathsf{value}$ $:\mathsf{Result}\times\mathsf{Context}\times\mathsf{Expr}\xrightarrow{}\mathsf{CValue.value}$ (non-pointer expression values)
@@
Figure 1: The interface of the backend analysis used by the proposed analysis of data races. The function $\mathsf{analyse\_thr}$ performs the analysis of a thread as a sequential program given by its entry point, starting from the given initial state. The other functions answer queries based on the results computed by $\mathsf{analyse\_thr}$ (specific for different backends). Note: the signatures do not include statements as those are already included in contexts.
@@ 4 Backend Analysis
Our data-race analysis relies on a backend analysis of sequential program behaviour to get answers for several kinds of queries. Those include function pointer resolution (to determine which functions are to be analysed when and to determine thread entry points passed to thread create functions), getting points-to information (to determine targets of (un)lock operations, to perform escape analysis, and, most importantly, to determine on which addresses to look for data races), getting information on memory accesses, and characterizing values of integer variables (to determine whether a $\mathsf{trylock}$ operation succeeds in the given branch of code).
To offer a selection of multiple backend analyses differing in their scalability, precision, and ease of use, we have implemented several such analyses. The interface a backend needs to implement is shown in Fig. 1. The function $\mathsf{analyse\_thr}$ runs the analysis of a thread as a sequential program starting from the provided state of global variables and the thread’s argument. Based on the results, the different queries are then answered.
@@
This backend is based on EVA [5], a value analysis plugin of Frama-C based on abstract interpretation. EVA provides a combination of multiple abstract domains and an API to answer all the needed queries, and our backend is just a thin wrapper over EVA, which implements caching, simplification, and answering some trivial queries on its own. The interface in Fig. 1 is in fact inspired by the API of EVA, and the other backends mimic it. In particular, the EVA’s API implements queries above call stacks which can be easily obtained from our contexts.
@@
This backend provides a lightweight analysis based only on syntactical information in the style of RacerX [13] or RacerD [3]. The implementation of $\mathsf{analyse}$ is simply an empty function and the functions $\mathsf{state}$ and $\mathsf{value}$ always return the top value. The function $\mathsf{value\_ptr}$ searches the expression and extracts bases of the expected type. This is inspired by the insight that in some programs only global locks are accessed using $\mathtt{lock(\&m)}$ and $\mathtt{unlock(\&m)}$ and it is thus sufficient to simply extract the base of $\mathtt{m}$. The function $\mathsf{memory\_accesses}$ is implemented similarly with a more detailed analysis of an expression to determine whether a base is being read or written to it. Finally, function pointers are over-approximated by taking each function assigned in the program. In the case of thread entry points, this can be further refined by taking only functions matching the thread signature (i.e., $\mathtt{void*(*fn)(void*arg)}$ for $\mathsf{pthreads}$).
@@
The aim of the third backend is to provide a middle ground between the precision of the EVA and the scalability of the syntactic backend. It relies on another Frama-C plugin called Alias [4] which performs a may-alias analysis using a variant of the classical Steensgaard’s algorithm [25]. For queries not related to pointer expressions ($\mathsf{state}$ and $\mathsf{value}$), the top values are returned. For the other queries, since Alias does not have API to answer our queries directly, we take the result of the syntactic backend and replace each base by a canonical representative of its alias equivalence class. In particular, for memory accesses, we ensure that a selected canonical address for an alias class is vulnerable to data races if such an address exists (e.g. because its base corresponds to a global variable or a local variable that escaped; see Section 7 for more details). Alias also does not allow to start with a user-supplied set of aliases (as expects the signature of our $\mathsf{analyse\_thr}$ function). We therefore need to perform an additional saturation of aliases set for all queries, as demonstrated by the example in the next paragraph. The Alias plugin is still relatively new (it lacks support, e.g., for the dynamic memory allocation) and thus our backend over it is also experimental.
Example.
@@
[obeytabs=true,tabsize=1]c void *t1(void *arg1) int *x = (int *)arg1; (*x)++;
void *t2(void *arg2) int *y = (int *)arg2; (*y)++;
int main() int data = 0; create(t1, &data); create(t2, data);
Figure 2: An example of a program for which syntactic backend fail to find a data race.
The advantage of the Alias backend over the syntactic backend is demonstrated in Fig. 2. The program contains a data race on an address of the variable $\mathtt{data}$ that escapes its scope in the $\mathtt{main}$ function by passing it as an argument to the threads $\mathtt{t1}$ and $\mathtt{t2}$. Both threads share the same code, which creates a local alias for this address. This is a very common pattern when thread arguments are used – since the type of the argument is generic $\mathtt{void*}$, it needs to be casted to its actual type. The address is then written to by both threads using their local aliases. While the syntactic backend would not report any relevant access for data race detection, the Alias backend would compute aliases for $\mathtt{x}$/$\mathtt{y}$ as $\{\mathtt{x},\mathtt{arg1}\}$/$\{\mathtt{y},\mathtt{arg2}\}$, respectively. Our backend then performs a saturation of both sets by adding aliases $\{\mathtt{arg1},\mathtt{\&data}\}$ and $\{\mathtt{arg2},\mathtt{\&data}\}$, resp. Finally, &$\mathtt{\&data}$ will be selected as a canonical alias base vulnerable to a data race.

### v2

@@ Abstract
We present RacerF, a novel static analyser for thread-modular data race detection. The approach behind RacerF exploits static analysis of sequential program behaviour whose results are generalised for multi-threaded programs using a combination of lightweight under- and over-approximating methods. The tool is implemented as a plugin of the Frama-C platform and can leverage several analysis backends, most notably the Frama-C’s abstract interpreter EVA. Although our methods are mostly heuristic without providing formal guarantees, our experimental evaluation shows that even for intricate programs, RacerF can provide very precise results competitive with more heavyweight approaches while being faster than them.
keywords:
concurrency, data race detection, static analysis
@@ 1 Introduction
A lot of work has also been invested in designing various static as well as dynamic methods of finding data races and/or proving their absence. However, the complex behaviour of concurrent programs reflects in that such methods do often have problems with scalability or precision. In this paper, we present a novel tool RacerF for thread-modular data race detection with the aim of providing a suitable balance between the scalability and precision of the analysis. We then also present our experience from running the tool and comparing it with several other data race detectors.
@@
The approach implemented in RacerF relies on leveraging results of static analysis of the sequential behaviour of program threads for finding data races in multi-threaded programs. We use two strategies of combining information resulting from the analysis of the sequential behaviour of program threads: one under-approximates and one over-approximates the concurrent behaviour of the threads running in parallel. We combine these strategies to obtain a more precise analysis such that the under-approximation is used to find data races while the over-approximation is used to show their absence.
@@
RacerF is implemented as a plugin of the Frama-C platform [10, 1] and is in particular designed for finding data races in C programs with POSIX threads (pthreads). RacerF can use several analysis backends implementing the underlying analysis of the sequential behaviour of program threads and differing in their precision and scalability. One of them is built on top of the Frama-C’s value analysis plugin EVA based on abstract interpretation [6], one is purely syntactic, and the last one is based on Alias, Frama-C’s plugin for may-alias and points-to analysis [5].
@@
We have performed experiments with RacerF on a number of benchmarks coming both from the SV-COMP competition in software verification, which includes both synthetic benchmarks as well as benchmarks derived from real-world code, as well as on a number of student projects in an advanced course of operating systems. Although the approach behind RacerF is mostly heuristic without providing formal guarantees, our experimental evaluation shows that even for intricate programs, RacerF can provide very precise results competitive with more heavy-weight approaches while being faster than them.
1.1 Approach Overview
Figure 1: An overview of the workflow of RacerF. The phases with under- and over-approximation guarantees are marked by blue and green colour, respectively. The phases designed without guarantees are marked by red colour. The solid lines represent the workflow of individual analysis phases. The dashed lines represent queries on the results of other phases.
An overview of the workflow of RacerF is illustrated in Fig. 1. The core component is a backend analyser which answers several kinds of queries based on analysing individual threads as sequential programs (the general interface and several different backend analysers are described in Section 4). The thread analysis (Section 5) incrementally discovers new program threads and analyses them using the chosen backend analyser. Since the discovery of a new thread may add additional information, the process of thread discovery and their analysis using the backend analyser is repeated until a fixpoint is reached. Then, two helper analyses are run to deal with checking whether two memory accesses may happen in parallel – a lockset analysis checks which accesses are guarded by which locks and an active-threads analysis checks when threads are created and joined (Section 6). Those analyses query results of the backend analysis, mainly for possible values of parameters of locking and threading functions. Finally, data race detection (Section 7) is performed. This phase checks memory accesses (using information provided by the backend analyser) and tracks whether they need to be protected and, if so, whether they indeed are. The detected issues are then classified into may- and must-data-races. Must-data-races are reported to the user, while an absence of may-data-races is used to claim program race-free.
Fig. 1 also depicts the approximation guarantees provided by the individual phases. Note that the combination of under- and over-approximating methods may seem problematic from the point of view of what can be guaranteed to hold for the produced results. However, we do not aim at providing guarantees in the sense of formal verification. Our aim is to achieve high efficiency and precision in practice even though some false alarms may be missed or some bugs not identified.
@@ 2 Related Work
RacerX [14] is a static analyser for C based primarily on a lockset analysis. The tool has been successfully used to find bugs in the kernels of several operating systems. Compared to our approach, RacerX uses a very simple pointer analysis for the representation of locks only. Further, RacerX uses many heuristics in individual analysis phases to deal with analysis mistakes while RacerF tries to achieve precision using careful ranking. RacerX is unfortunately not available for an experimental comparison. RacerD [4] is a compositional analyser implemented in the Infer framework. It provides scalability by analysing each function only once. Unlike us, RacerD targets Java programs with structural locking, assuming that all locking operations are balanced. Both RacerX and RacerD focus on under-approximation only. O2 [25] is another static checker building on the notion of origins unifying threads and events. The approach was implemented for C, C++, and Java/Android; and used to find many previously unknown bugs in real-world code. However, as we show in our experiments, it can still miss many data races in smaller but real-world programs.
@@
On the other side, there are tools focused on soundness, rather than data race detection. Goblint [31] and Astrée [23, 26] rely on a thread-modular abstract interpretation to soundly over-approximate thread behaviour. The work [31] further proposes a way to handle common locking schemes in Linux driver devices. In our approach, we are not interested in the analysis of multithreaded programs in general and we can consequently afford more drastic approximations that seem to be sufficient for data race detection. We also focus on under-approximation to detect bugs.
@@
As an alternative to static data race detection, various approaches based on testing and dynamic analysis can be used. A big challenge for these approaches is the fact that errors like data races can manifest in very rare behaviours only, which are difficult to catch, even through many repeated test runs. Coverage of such rare behaviours can be improved using approaches such as systematic testing [32] or noise-based testing [12, 16]. Another way is to use extrapolating dynamic checkers, such as [29, 13, 18], which can report warnings about possible errors even if those are not seen in the testing runs, based on spotting their symptoms. However, even though such checkers have proven quite useful in practice, they can, of course, still miss rarely occurring errors. Moreover, monitoring a run of large software through such checkers may be quite expensive too.
@@
Approaches based on model checking, i.e., exhaustive state-space exploration, can guarantee the discovery of all potentially present errors – either in general or at least up to some bound, often given in the number of context switches. However, so far, the scalability of these techniques is not sufficient to handle large industrial code, even when combined with methods such as sequentialisation [24, 27] or advanced techniques based on automata [11, 22] or SMT encodings [19, 20, 15]. Indeed, this shows up in our experiments too.
@@ 3 Preliminaries
We now introduce several basic notions that we use when describing the analysed programs and the analyses proposed.
@@
The CValue domain and memory addresses.
Throughout the paper, we frequently refer to the $\mathsf{CValue}$ abstract domain [6, 10] implemented in Frama-C and used by the abstract interpretation of EVA. We are primarily interested in the representation of memory addresses. A memory address is represented as a pair $(b,o)$ where $b$ is a base address (in the following referred to simply as a base) and $o$ is an offset. Each variable (global, local, and also formal) has its associated base address. Dynamic allocations are modelled by creating so-called dynamic bases which come in two flavours – a strong dynamic base corresponds to the result of precisely one call to an allocation function; on the other hand, weak memory bases may represent multiple allocated addresses (typically, they are the consequence of allocations in loops). Offsets are represented by a dedicated integer domain, the details of which are not important for this paper. We use $\mathsf{CValue.val}$ for the domain representing possible values of a single variable and $\mathsf{CValue.state}$ for the domain representing program state, i.e., possible values for multiple variables. We use $\mathsf{Addresses}$ to denote the aforementioned set of memory addresses. In the rest of the paper, the operations join and widening as well as top values refer to those on the $\mathsf{CValue}$ domain.
@@
We call an analysis context a triple $(thread,calls,stmt)$ consisting of a thread, a sequence of the latest $n$ calls (each of them formed by a call site and the function called), and a program statement. Our data race analysis relies on various underlying analyses presented below to compute an abstract representation of thread states reachable by performing the given statement for each analysis context. Notice that a context holds no information about how the thread was created nor how its execution was interleaved with other threads. The empty call sequence is denoted by $\varepsilon$. The number $n$ is a parameter of the analysis set to 1 by default to handle simple lock/unlock wrapper functions. We use $\mathsf{Context}$ to denote the set of all possible contexts.
$\mathsf{analyse\_thread}$ $:\mathsf{Function}\times\mathsf{CValue.state}\xrightarrow{}\mathsf{Result}$ (sequential analysis of a thread)
@@
$\mathsf{value}$ $:\mathsf{Result}\times\mathsf{Context}\times\mathsf{Expr}\xrightarrow{}\mathsf{CValue.val}$ (non-pointer expression values)
@@
Figure 2: The interface of the backend analysis used by the proposed analysis of data races. The function $\mathsf{analyse\_thread}$ performs the analysis of a thread as a sequential program given by its entry point, starting from the given initial state. The other functions answer queries based on the results computed by $\mathsf{analyse\_thread}$ (specific for different backends). Note that the signatures do not include statements as those are already included in contexts.
@@ 4 Backend Analysis
Our data race analysis relies on a backend analysis of sequential program behaviour to get answers for several kinds of queries. Those include function pointer resolution (to determine which functions are to be analysed when and to determine thread entry points passed to thread create functions), getting points-to information (to determine targets of (un)lock operations, to perform escape analysis, and, most importantly, to determine on which addresses to look for data races), getting information on memory accesses, and characterizing values of integer variables (to determine whether offsets of two accesses overlap, and to check whether a $\mathsf{trylock}$ operation succeeds in the given branch of code).
To offer a selection of multiple backend analyses differing in their scalability, precision, and ease of use, we have implemented several such analyses. The interface a backend needs to implement is shown in Fig. 2. The function $\mathsf{analyse\_thread}$ runs the analysis of a thread as a sequential program starting from the provided state of global variables and the thread’s argument. Based on the results, the different queries are then answered.
@@
This backend is based on EVA [6], a value analysis plugin of Frama-C based on abstract interpretation. EVA provides a combination of multiple abstract domains and an API to answer all the needed queries, and our backend is just a thin wrapper over EVA, which implements caching, simplification, and answering some trivial queries on its own. The interface in Fig. 2 is in fact inspired by the API of EVA, and the other backends mimic it. In particular, the EVA’s API implements queries for call stacks which can be easily obtained from our contexts.
@@
This backend provides a lightweight analysis based only on syntactical information in the style of RacerX [14] or RacerD [4]. The implementation of $\mathsf{analyse\_thread}$ is simply an empty function and the functions $\mathsf{state}$ and $\mathsf{value}$ always return the top value. The function $\mathsf{value\_ptr}$ searches the expression and extracts bases of the expected type. This is inspired by the insight that in some programs only global locks are accessed using $\mathtt{lock(\&m)}$ and $\mathtt{unlock(\&m)}$ (where $\mathtt{m}$ is a global variable), and it is thus sufficient to simply extract the base of $\mathtt{m}$. The function $\mathsf{memory\_accesses}$ is implemented similarly with a more detailed analysis of an expression to determine whether a base is being read or written to it. Finally, function pointers are over-approximated by taking each function assigned in the program. In the case of thread entry points, this can be further refined by taking only functions matching the thread signature (i.e., $\mathtt{void*(*fn)(void*arg)}$ for $\mathsf{pthreads}$).
@@
The aim of the third backend is to provide a middle ground between the precision of the EVA and the scalability of the syntactic backend. It relies on another Frama-C plugin called Alias [5] which performs a may-alias analysis using a variant of the classical Steensgaard’s algorithm [30]. For queries not related to pointer expressions ($\mathsf{state}$ and $\mathsf{value}$), the top values are returned. For the other queries, since Alias does not have an API to answer our queries directly, we take the result of the syntactic backend and replace each base by a canonical representative of its alias equivalence class. In particular, for memory accesses, we ensure that a selected canonical address for an alias class is vulnerable to data races if such an address exists (e.g. because its base corresponds to a global variable or a local variable that escaped; see Section 7 for more details). Alias also does not allow starting with a user-supplied set of aliases (as expects the signature of our $\mathsf{analyse\_thread}$ function). We therefore need to perform an additional saturation of the aliases set for all queries, as demonstrated by the example in the next paragraph. The Alias plugin is still relatively new (it lacks support, e.g., for the dynamic memory allocation) and thus our backend over it is also experimental.
@@
[numbersep=-6pt, obeytabs=true,tabsize=1,linenos]c void *t1(void *arg1) int *x = (int *)arg1; (*x)++;
void *t2(void *arg2) int *y = (int *)arg2; (*y)++; int main() int data = 0;
create(t1, &data); create(t2, data); a
a
Figure 3: An example of a program for which the Syntactic backend fails to find a data race.
Example 4.1.
The advantage of the Alias backend over the syntactic backend is demonstrated in Fig. 3. The program contains a data race on an address of the variable $\mathtt{data}$ that escapes its scope in the $\mathtt{main}$ function by passing it as an argument to the threads $\mathtt{t1}$ and $\mathtt{t2}$. Both threads share the same code, which creates a local alias for this address. This is a very common pattern when thread arguments are used – since the type of the argument is generic $\mathtt{void*}$, it needs to be casted to its actual type. The address is then written to by both threads using their local aliases. While the syntactic backend would not report any relevant access for data race detection, the Alias backend would compute aliases for $\mathtt{x}$ and $\mathtt{y}$ as $\{\mathtt{x},\mathtt{arg1}\}$ and $\{\mathtt{y},\mathtt{arg2}\}$, respectively. Our backend then performs a saturation of both sets by adding aliases $\{\mathtt{arg1},\mathtt{\&data}\}$ and $\{\mathtt{arg2},\mathtt{\&data}\}$, respectively. Finally, &$\mathtt{\&data}$ will be selected as a canonical base, because it is vulnerable to a data race.